The series

1/3+ 1/7+ 1/11 + 1/15+.....

is a convergent series.

True or false with full explanation


1
Expert's answer
2021-03-30T07:30:43-0400

The series

1/3+ 1/7+ 1/11 + 1/15+.....

is a convergent series.

True or false with full explanation

Solution:

"\frac13+\frac17+\frac{1}{11}+\frac{1}{15}+...=\displaystyle\sum_{n=1}^\infty\frac{1}{4n-1}=\\\frac14\displaystyle\sum_{n=1}^\infty\frac{1}{n-\frac14}=\frac14\displaystyle\sum_{n=1}^\infty a_n"

Let's apply direct comparison test:

"a_n=\displaystyle\frac{1}{n-\frac14}>\frac{1}{n}"

"\displaystyle\sum_{n=1}^\infty\frac{1}{n}" is the harmonic series that is the divergent infinite series.

If "\displaystyle\sum_{n=1}^\infty\frac{1}{n}" is a divergent series and "a_n>\frac{1}{n}" then "\displaystyle\sum_{n=1}^\infty a_n" is also a divergent series.

Since

"\displaystyle\sum_{n=1}^\infty a_n" is a divergent series

than

"\frac14\displaystyle\sum_{n=1}^\infty a_n=\frac13+\frac17+\frac{1}{11}+\frac{1}{15}+..." is also a divergent series.

Answer: false.